# Comparing Left and Right Quotient Sets in Groups

### v2

@@
Abstract.
For a finite subset $A$ of a group $G$, we define the right quotient set and the left quotient set of $A$, respectively, as
$AA^{-1}$ $~\coloneqq~\{a_{1}a_{2}^{-1}:a_{1},a_{2}\in A\},$
$A^{-1}A$ $~\coloneqq~\{a_{1}^{-1}a_{2}:a_{1},a_{2}\in A\}.$
While the right and left quotient sets are equal if $G$ is abelian, subtleties arise when $G$ is a nonabelian group, where the cardinality difference $|AA^{-1}|-|A^{-1}A|$ may be take on arbitrarily large values. Using the results of Martin and O’Bryant on the cardinality differences of sum sets and difference sets in $\mathbb{Z}$, we prove in the infinite dihedral group, $D_{\infty}\cong\mathbb{Z}\rtimes\mathbb{Z}/2\mathbb{Z}$, every integer difference is achievable. Further, we prove that in $F_{2}$, the free group on $2$ generators, an integer difference is achievable if and only if that integer is even, and we explicitly construct subsets of $F_{2}$ that achieve every even integer. We further determine the minimum cardinality of $A\subset G$ so that the difference between the cardinalities of the left and right quotient sets is nonzero, depending on the existence of order $2$ elements in $G$. To prove these results, we construct difference graphs $D_{A}$ and $D_{A^{-1}}$ which encode equality, respectively, in the right and left quotient sets. We observe a bijection from edges in $D_{A}$ to edges in $D_{A^{-1}}$ and count connected components in order to obtain our results on cardinality differences $|AA^{-1}|-|A^{-1}A|$.
This research was supported by the National Science Foundation, under NSF Grant DMS2241623, Amherst College, Princeton University, the University of Wisconsin, the University of Michigan, Williams College, and the Finnerty Fund. We thank the participants of the 2025 Integers Conference for conversations indirectly inspiring this work.
1. Introduction
Note: This paper is dedicated with thanks to Carl Pomerance and Mel Nathanson. The genesis for this work came from conversations the fifth named author had at the Integers Conference in Georgia in 2025 on Results in Additive & Elementary Number Theory Inspired by Carl and Mel. His presentation described how much of his mentoring of students has been influenced by each, in particular problems in the orbit of MSTD sets (from Mel) and multiplicative and quotient structures (from Carl). This led to springboard problems for the SMALL 2025 REU, which led to the work below.
1.1. Background
@@
$A-A$ $~\coloneqq~\{a_{1}-a_{2}:a_{1},a_{2}\in A\}.$
A natural comparison arises between the cardinalities of the sum and difference sets. Our set $A$ is said to be sum dominated or MSTD (more sums than differences) if |A+A|>|AA||A+A|>|A-A| and difference dominated (MDTS) if |AA|>|A+A||A-A|>|A+A|. One might expect that the difference set would have a greater cardinality as addition is commutative in $\mathbb{Z}$ while subtraction is not. In particular, if we consider a pair of distinct elements $a_{1},a_{2}\in A$, $a_{1}-a_{2}$ and $a_{2}-a_{1}$ are distinct elements in the difference set, while $a_{1}+a_{2}=a_{2}+a_{1}$ is a single element in the sumset. However, it is possible to have a set that is sum dominated. The earliest example of a MSTD set was discovered by Conway in 1960s: $\{0,2,3,4,7,11,12,14\}$. Surprisingly, in contrast to the intuition that MSTD sets should form a vanishing proportion of subsets of $[N]$ as $N$ grows large, in 2006 Martin and O’Bryant [MO06] proved that the proportion of subsets of $N]\subseteq\mathbb{Z}$ which are MSTD does not vanish as $N\to\infty$. Since then, extensive research has expanded the classical MSTD problem to various settings, including higher dimensions and various families [KM22, CLMS20, CLMS19].
One natural extension of the study of MSTD sets amongst subsets of $\mathbb{Z}$ is to ask this question for groups in general. We primarily focus on the free group $F_{2}$ on two generators and the infinite dihedral group $D_{\infty}\cong\mathbb{Z}\rtimes\mathbb{Z}/2\mathbb{Z}$.
@@
Following previous work by [ACJ+22, MV14, Nat07, Zha10], we look at sumsets and difference sets in general groups. Let $A\subset G$ be a finite subset of a group $G$ equipped with the operation $\times:(x,y)\mapsto xy$ and the inverse map $^{-1}:x\mapsto x^{-1}$. Although some authors use the terminology sumset and difference set in this context [ACJ+22, MV14, Nat07, Zha10], we will refer to the product set and the quotient set of $A$, which are defined, respectively as
@@
$AA^{-1}$ $~\coloneqq~\{xy^{-1}:x,y\in A\}.$
More generally, let $G$ be any group. Let $A=\left\{a_{1},\dots,a_{n}\right\}\subseteq G$. Define $A^{-1}\coloneqq\left\{a_{1}^{-1},\dots,a_{n}^{-1}\right\}$. We consider the right and left quotient sets of $A$, defined, respectively, by
@@
$A^{-1}A~\coloneqq~\left\{a_{i}^{-1}\cdot a_{j}:a_{i},a_{j}\in A\right\}.$
In the classical MSTD problem, one natural area of study has been the cardinality of the smallest MSTD set. Hegarty showed that the smallest such example in $\mathbb{Z}$ has cardinality $8$, and is unique up to affine translation [Heg07]. Another natural question asks for the set of values |A+A||AA||A+A|-|A-A| can attain across all $A\subset[N]$. Martin and O’Bryant proved, over finite subsets of $\mathbb{Z}$, that this difference achieves all integers [MO06].
Inspired by this, we consider similar questions about left and right quotient sets. First note that the problem of constructing a set $A$ where $|AA^{-1}|>|A^{-1}A|$ is the same as the problem of $|AA^{-1}|<|A^{-1}A|$ (by replacing $A$ with $A^{-1}$). Therefore, a more natural question to ask is the smallest $A$ where the left and right quotient sets are not equal. Moreover, inspired by [MO06], we give results on what values $|AA^{-1}|-|A^{-1}A|$ can take as $A$ ranges over finite subsets of a group $G$.
Example $1.$
While there has been some work done generalizing in a abelian group setting [MV14, Nat07, Zha11], our questions are only relevant for non-abelian groups. Indeed, if $G$ is abelian, then all $A\subseteq G$ have $AA^{-1}=A^{-1}A$, so the only possible value of $|AA^{-1}|-|A^{-1}A|$ is 0.
1.2. Notation and Main Results
@@
Let $G$ be a group with no elements of order 22.111While we use the language of “a group with no elements of order 22”, we remark that Tao refers to such a group as a “a $2$-torsion-free group” in [Tao08]. Let $A\subseteq G$ be a finite subset. Then $|AA^{-1}|-|A^{-1}A|$ is even.
@@
By [MO06, Theorem 4], the latter difference ranges over every integer.
@@
A brute force search on groups of small order show that the bound $|A|\geq 4$ is sharp (see 2.2). The main tool we use to prove this result is a graph associated to the left (and right) quotient sets, which we call the difference graph. The cardinalities of $AA^{-1}$ and $A^{-1}A$ can be interpreted as the number of connected components on these graphs. Making use a bijection of edges between these graphs, we perform an argument based on the properties of this graph to prove that when $|A|\leq 3$ (resp. $|A|\leq 4$ when $A$ has no elements of order $2$), the number of connected components does not change under the bijection of edges.
2. Left vs. Right Quotient Sets
2.1. Graph Construction
To prove our results, we define the difference graph $D_{A}$ of a finite subset $A\coloneqq\{a_{1},a_{2},...,a_{n}\}\subseteq G$. The graph $D_{A}=(V,E)$ is defined as follows.
  1. (1)
@@ 2.1. Graph Construction
      If $G$ has no elements of order $2$, $[(i,j),(j,i)]$, an edge connecting to its symmetric pair, provided that $j\neq i$.
  3. (3)
@@
$a_{i}a_{j}^{-1}~=~a_{i}a_{k}^{-1}\implies a_{j}~=~a_{k}.$
(2c) Let $(i,j),(j,i)]\in E(D_{A})$ where $i\neq j$.
$a_{i}a_{j}^{-1}~=~a_{j}a_{i}^{-1}~=~(a_{i}a_{j}^{-1})^{-1}\implies a_{i}a_{j}^{-1}~=~e\implies a_{i}~=~a_{j},$
contradicting our asusmption that $i\neq j$.
@@
In light of Property (4), we will refer to a connected component $C$ in $D_{A}$ consisting of $k$ elements simply by indicating its vertices $C=(a_{1},b_{1})(a_{2},b_{2})\cdots(a_{n},b_{n})$.
@@
A priori, this map is only well-defined if we consider edges as directed (the reverse edge gets mapped to the transpose of the original edge) and allow loops (they get mapped to the diagonal). However, since $T$ is an automorphism of $D_{A}$, we may take $D_{A}$ to be undirected. This is a well-motivated map: its existence already tells us that the additive energies $\Lambda(A,A^{-1})~\coloneqq~\{\#(a_{1},a_{2},a_{3},a_{4})\in A^{4}:a_{1}a_{2}^{-1}=a_{3}a_{4}^{-1}\}$ and $\Lambda(A^{-1},A)~\coloneqq~\{\#(a_{1},a_{2},a_{3},a_{4})\in A^{4}:a_{1}^{-1}a_{2}=a_{3}^{-1}a_{4}\}$ are equal in non-commutative groups (see [Tao08]).
@@
When $n=5$, we consider the following example.
@@
2.2. Possible Differences
The natural question that arises is: what are the possible differences between the cardinalities of the right and left quotient sets of $A\subset G$? In 1.2 we gave a construction on the infinite dihedral group $D_{\infty}$ demonstrating the difference $|AA^{-1}|-|A^{-1}A|$ achieves every possible $n\in\mathbb{Z}$. Given the restriction that there are no elements of order 2 in our group, we can conclude that $|AA^{-1}|-|A^{-1}A|$ is even. The proof of Theorem 1.1 follows.
@@ 2.2. Possible Differences
It suffices to show that $c(D_{A})-c(D_{A^{-1}})$ is even. We claim that $c(D_{A})$ is odd.
We divide $\mathcal{C}(D_{A})$ into two disjoint classes.
  1. (1)
@@
Denote these sets, respectively, by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Note that $\left\lvert\mathcal{C}_{2}\right\rvert$ is even as each component comes in pairs.
We claim that $\mathcal{C}_{1}$ contains only the diagonal (the diagonal is a connected component by 2.1 (3, 2a)). Indeed, if any other component $C$ belongs to $\mathcal{C}_{1}$, then there exists a vertex $(i,j)$ with $i\neq j$ such that $(i,j),(j,i)]\in C$, which contradicts 2.1 (2c). Therefore, there is exactly one component of $\mathcal{C}_{1}$ while the rest of the connected components come in symmetric pairs, so $|\mathcal{C}_{1}|$ is odd. Applying the same reasoning to $c(D_{A^{-1}})$ shows it is odd, so $c(D_{A})-c(D_{A^{-1}})$ is even. ∎
@@
Any set $A\subseteq\left\langle a\right\rangle\cong\mathbb{Z}$ with $a\in F_{2}$ yields the $n=0$ case. By replacing $A$ with $A^{-1}$, it suffices to prove the result for positive $n$. So we construct a family of sets $A_{n}\subseteq F_{m}$ for some $m\geq 2$ and then compose it with the embedding (2.4). The following set (described above) in $F_{3}=F(\left\{x,y,z\right\})$ for $n=1$
@@
This implies [^2] For $i\neq j$, the sets $A_{n}^{(i)}$ and $A_{n}^{(j)}$ are supported on disjoint sets of generators in $F_{3k}$, meaning any product of the form $ab^{-1}$ or $a^{-1}b$ where $a\in A_{n}^{(i)}$ and $b\in A_{n}^{(j)}$ produces a word involving letters from distinct alphabets. Thus all such products are distinct. However, if $i=j$, the words produced come from the same alphabet, which doesn’t generate distinct products.
@@
Note that $|C_{k}|=2k$ and thus let $n=2k$ . Next, we will compute the size of the right quotient set:
@@
The following remark is mentioned in [Tao08]: if $H\subseteq G$ is a finite, nonempty subset of some group $G$ and $g$ is an element of $G$ not in the normalizer of $H$, then the set $A\coloneqq Hg\cup H$ has $A^{-1}A$ about the same size as $H$, but $AA^{-1}$ can be possibly large. Our example in $F_{2}$ is this construction with $H=\left\{x,x^{2},\dots,x^{k}\right\}$ and $g=y$.
2.3. Cardinality of AA
The set $A$ given in (2.1) is extremal in the sense that it is the smallest such set such that $|AA^{-1}|-|A^{-1}A|$ in $F_{2}$. In fact, this holds for all groups with no elements of order $2$, which we prove in Theorem 1.3.

### v3

# INTEGERS 26 (2026)

@@
Abstract
For a finite subset $A$ of a group $G$, we define the right quotient set and the left quotient set of $A$ as
$AA^{-1}$ $\ ~:=~\ \{a_{1}a_{2}^{-1}:a_{1},a_{2}\in A\},$
$A^{-1}A$ $\ ~:=~\ \{a_{1}^{-1}a_{2}:a_{1},a_{2}\in A\}$
respectively. While the right and left quotient sets are equal if $G$ is abelian, subtleties arise when $G$ is a non-abelian group, where the cardinality difference $|AA^{-1}|-|A^{-1}A|$ may take on arbitrarily large values. Using the results of Martin and O’Bryant on the cardinality differences of sum sets and difference sets in $\mathbb{Z}$, we prove that in the infinite dihedral group, $D_{\infty}\cong\mathbb{Z}\rtimes\mathbb{Z}/2\mathbb{Z}$, every integer difference is achievable. Further, we prove that in $F_{2}$, the free group on $2$ generators, an integer difference is achievable if and only if that integer is even, and we explicitly construct subsets of $F_{2}$ that achieve every even integer. We further determine the minimum cardinality of $A\subset G$ so that the difference between the cardinalities of the left and right quotient sets is nonzero, depending on the existence of order $2$ elements in $G$. To prove these results, we construct difference graphs $D_{A}$ and $D_{A^{-1}}$ which encode equality in the right and left quotient sets respectively. We observe a bijection from edges in $D_{A}$ to edges in $D_{A^{-1}}$ and count connected components in order to obtain our results on cardinality differences $|AA^{-1}|-|A^{-1}A|$.
§1  Introduction
This paper is dedicated with thanks to Carl Pomerance and Mel Nathanson. The genesis for this work came from conversations the fifth named author had at the Integers Conference in Georgia in 2025 on Results in Additive & Elementary Number Theory Inspired by Carl and Mel. His presentation described how much of his mentoring of students has been influenced by each, in particular problems in the orbit of MSTD sets (from Mel) and multiplicative and quotient structures (from Carl). This led to springboard problems for the SMALL 2025 REU, which led to the work below.
1.1 Background
@@
$A-A$ $\ ~\coloneqq~\ \{a_{1}-a_{2}:a_{1},a_{2}\in A\}$
respectively. A natural comparison arises between the cardinalities of the sum and difference sets. Our set $A$ is said to be sum dominated or MSTD (more sums than differences) if |A+A|>|AA||A+A|>|A-A| and difference dominated (MDTS) if |AA|>|A+A||A-A|>|A+A|. One might expect that the difference set would have a greater cardinality as addition is commutative in $\mathbb{Z}$ while subtraction is not. In particular, if we consider a pair of distinct elements $a_{1},a_{2}\in A$, $a_{1}-a_{2}$ and $a_{2}-a_{1}$ are distinct elements in the difference set, while $a_{1}+a_{2}=a_{2}+a_{1}$ is a single element in the sumset. However, it is possible to have a set that is sum dominated. The earliest examples of MSTD sets were discovered in the 1960s by Conway ($\{0,2,3,4,7,11,12,14\}$). Surprisingly, against the intuition that MSTD sets should form a vanishing proportion of subsets of $[N]$ as $N$ grows large, in 2006 Martin and O’Bryant [9] proved that the proportion of subsets of $N]\subseteq\mathbb{Z}$ which are MSTD does not vanish as $N\to\infty$. Since then, extensive research has expanded the classical MSTD problem to various settings, including higher dimensions and various families [6, 2, 3].
A natural extension of the study of MSTD sets among subsets of $\mathbb{Z}$ is to ask the question for general groups. This has been investigated in previous work by [1, 10, 12, 15]. In this paper, the main groups of interest are the free group $F_{2}$ on two generators and the infinite dihedral group $D_{\infty}\cong\mathbb{Z}\rtimes\mathbb{Z}/2\mathbb{Z}$.
@@
We now define sumsets and difference sets in general groups. Let $A\subset G$ be a finite subset of a group $G$ equipped with the operation $\times:(x,y)\mapsto xy$ and the inverse map $^{-1}:x\mapsto x^{-1}$. Although some authors use the terminology sumset and difference set in this context [1, 10, 12, 15], we refer to the product set and the quotient set of $A$, which are defined as
@@
$AA^{-1}$ $\ ~\coloneqq~\ \{xy^{-1}:x,y\in A\}$
respectively.
More generally, let $G$ be any group. Let $A=\left\{a_{1},\dots,a_{n}\right\}\subseteq G$. Define $A^{-1}\coloneqq\left\{a_{1}^{-1},\dots,a_{n}^{-1}\right\}$. We consider the right quotient set and left quotient set of $A$, defined by
@@
$A^{-1}A\ ~\coloneqq~\ \left\{a_{i}^{-1}\cdot a_{j}:a_{i},a_{j}\in A\right\}$
respectively. In the classical MSTD problem, one natural area of study has been the cardinality of the smallest MSTD set. Hegarty showed that the smallest such example in $\mathbb{Z}$ has cardinality $8$, and is unique up to affine translation [4]. Another natural question asks for the set of values that |A+A||AA||A+A|-|A-A| can attain across all $A\subset[N]$. Martin and O’Bryant proved, over finite subsets of $\mathbb{Z}$, that this difference achieves all integers [9].
Inspired by this, we consider similar questions about left and right quotient sets. First note that the problem of constructing a set $A$ where $|AA^{-1}|>|A^{-1}A|$ is the same as the problem of $|AA^{-1}|<|A^{-1}A|$ (by replacing $A$ with $A^{-1}$). Therefore, a more natural question to ask is “What is the smallest $A$ where the left and right quotient sets are not equal?” Inspired by [9], we give results on what values $|AA^{-1}|-|A^{-1}A|$ can take as $A$ ranges over finite subsets of a group $G$.
There has been work generalizing MSTD questions to abelian groups [10, 12, 15]. However, our questions are only relevant for non-abelian groups since $AA^{-1}=A^{-1}A$ for $A$ a subset of an abelian group.
1.2 Notation and Main Results
@@
Let $G$ be a group with no elements of order 22.111While we use the language of “a group with no elements of order 22”, we remark that Tao refers to such a group as a “a $2$-torsion-free group” in [13]. Let $A\subseteq G$ be a finite subset. Then $|AA^{-1}|-|A^{-1}A|$ is even.
@@
By [9, Theorem 4], the latter difference ranges over every integer.
@@
A brute force search among groups of small order shows that the bound $|A|\geq 4$ is sharp (see Example 2.10). The main tool we use to prove this result is a graph associated to the left (and right) quotient sets, which we call the difference graph. The cardinalities of $AA^{-1}$ and $A^{-1}A$ can be interpreted as the number of connected components on these graphs. Making use of a bijection of edges between these graphs, we perform an argument based on the properties of this graph to prove that when $|A|\leq 3$ (resp. $|A|\leq 4$ when $A$ has no elements of order $2$), the number of connected components does not change under the bijection of edges.
§2  Left vs. Right Quotient Sets
2.1 Graph Construction
To prove our results, we define the difference graph $D_{A}$ of a finite subset $A\coloneqq\{a_{1},a_{2},...,a_{n}\}\subseteq G$. The graph $D_{A}=(V,E)$ is defined as follows:
  1. 1.
@@ 2.1 Graph Construction
      If $G$ has no elements of order $2$, then no edge connects a vertex $(i,j)$ to its symmetric pair $(j,i)$, provided that $j\neq i$.
  3. 3.
@@
$a_{i}a_{j}^{-1}\ ~=~\ a_{i}a_{k}^{-1},$
which implies $a_{j}~=~a_{k}$.
(2c) Let $(i,j),(j,i)]\in E(D_{A})$ where $i\neq j$. Then,
$a_{i}a_{j}^{-1}\ ~=~\ a_{j}a_{i}^{-1}\ ~=~\ (a_{i}a_{j}^{-1})^{-1},$
which implies
$(a_{i}a_{j}^{-1})^{2}\ ~=~\ e.$
Because $G$ has no elements of order $2$, so $a_{i}a_{j}^{-1}=e$ and $a_{i}~=~a_{j}$, contradicting our assumption that $i\neq j$.
@@
In light of Property (4) Lemma 2.1, we refer to a connected component $C$ in $D_{A}$ consisting of $k$ elements simply by indicating its vertices $C=(a_{1},b_{1})(a_{2},b_{2})\cdots(a_{k},b_{k})$.
@@
A priori, this map is only well-defined if we consider edges as directed (the reverse edge gets mapped to the transpose of the original edge) and allow loops (they get mapped to the diagonal). However, since $T$ is an automorphism of $D_{A}$, we may take $D_{A}$ to be undirected. This is a well-motivated map: its existence already tells us that the additive energies $\Lambda(A,A^{-1})~\coloneqq~\{\#(a_{1},a_{2},a_{3},a_{4})\in A^{4}:a_{1}a_{2}^{-1}=a_{3}a_{4}^{-1}\}$ and $\Lambda(A^{-1},A)~\coloneqq~\{\#(a_{1},a_{2},a_{3},a_{4})\in A^{4}:a_{1}^{-1}a_{2}=a_{3}^{-1}a_{4}\}$ are equal in non-commutative groups (see [13]).
@@
When $n=5$, consider the example in Figure 1.
@@
Figure 1: This example shows that the number of connected components in $D_{A}$ changes after applying the map $\phi$.
@@
2.2 Possible Differences
The natural question that arises is: what are the possible differences between the cardinalities of the right and left quotient sets of $A\subset G$? In Example 1.3 we gave a construction on the infinite dihedral group $D_{\infty}$ demonstrating the difference $|AA^{-1}|-|A^{-1}A|$ achieves every possible $n\in\mathbb{Z}$. Given the restriction that there are no elements of order 2 in our group, we can conclude that $|AA^{-1}|-|A^{-1}A|$ is even. The proof of Theorem 1.2 follows.
@@ 2.2 Possible Differences
It suffices to show that $c(D_{A})-c(D_{A^{-1}})$ is even. We claim that $c(D_{A})$ is odd. We divide $\mathcal{C}(D_{A})$ into two disjoint classes.
  1. 1.
@@
Denote these sets by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively. Note that $\left\lvert\mathcal{C}_{2}\right\rvert$ is even as each component comes in pairs.
We claim that $\mathcal{C}_{1}$ contains only the diagonal (the diagonal is a connected component by Lemma 2.1 (3, 2a)). Indeed, if any other component $C$ belongs to $\mathcal{C}_{1}$, then there exists a vertex $(i,j)$ with $i\neq j$ such that $(i,j),(j,i)]\in C$, which contradicts Lemma 2.1 (2c). Therefore, there is exactly one component of $\mathcal{C}_{1}$ while the rest of the connected components come in symmetric pairs, so $|\mathcal{C}_{1}|$ is odd. Applying the same reasoning to $c(D_{A^{-1}})$ shows it is odd, so $c(D_{A})-c(D_{A^{-1}})$ is even. ∎
@@
Any set $A\subseteq\left\langle a\right\rangle\cong\mathbb{Z}$ with $a\in F_{2}$ yields the $n=0$ case. By replacing $A$ with $A^{-1}$, it suffices to prove the result for positive $n$. So, we construct a family of sets $A_{n}\subseteq F_{m}$ for some $m\geq 2$ and then compose it with the embedding in Fact 2.5. The following set (described above) in $F_{3}=F(\left\{x,y,z\right\})$ for $n=1$
@@
This implies [^2] For $i\neq j$, the sets $A_{n}^{(i)}$ and $A_{n}^{(j)}$ are supported on disjoint sets of generators in $F_{3k}$, meaning any product of the form $ab^{-1}$ or $a^{-1}b$ where $a\in A_{n}^{(i)}$ and $b\in A_{n}^{(j)}$ produces a word involving letters from distinct alphabets. Thus, all such products are distinct. However, if $i=j$, the words produced come from the same alphabet, which do not generate distinct products.
@@
Note that $|C_{k}|=2k$ and thus let $n=2k$. Next, we compute the size of the right quotient set:
@@
The following remark is mentioned in [13, Remark 4.4.]. If $H\subseteq G$ is a finite subgroup of some group $G$ and $g$ is an element of $G$ not in the normalizer of $H$, then the set $A\coloneqq Hg\cup H$ has $A^{-1}A$ about the same size as $H$, but $AA^{-1}$ can be possibly large. If we slightly modify this construction, letting $H$ be a subset of $G$, then the proof of Proposition 2.6 is the case $H=\left\{x,x^{2},\dots,x^{k}\right\}$ and $g=y$. Indeed, $|A^{-1}A|=O(|H|)$, but $|AA^{-1}|=O(|H|^{2})$.
2.3 Cardinality of AA
The set $A$ given in (2.1) is extremal in the sense that it is a subset of $F_{2}$ with minimal cardinality such that $|AA^{-1}|\neq|A^{-1}A|$. In fact, it has minimal cardinality among subsets with $|AA^{-1}|\neq|A^{-1}A|$ in any group with no elements of order $2$ as a consequence of Theorem 1.5.